Two vectors A and B lie in a plane.Another Vector C lies outside this Plane.The Resultatnt A+B+C of these three Vectors.
(a)can be zero
(b)cannot be zero
(c)Lies in the plane of A and B
(d)Lies in the Plane of A and A+B

Grade:11

8 Answers

kartikay guleria
11 years ago

cant be zero ... 

Nimish Singh
11 years ago

Can you tell me the reason?

xyz xz
11 years ago

a vector is nullified if and only if another vector attached to its tail in opp. direction ofsame magnitude n SAME PLANE only.

 

A and B forms a resultant vector in AB plane and the resultant of A+B+C cant be zero

Abhishekh kumar sharma
10 years ago

cannot be zero cleary

ganapathy
7 years ago
Since the direction is not defined both the A and B vectors are added so it cant be zero this is the practical imagination of vectors
Ishan karki
6 years ago
It cannot be zero because it is mentioned that the three of the vector are not in same plane. Only two vectors are in same plane. So the resultant of three vectors can never be zero.
 
Anuj
5 years ago
As it is given that the vector A and B lies in the a plane,so their resultant must lies in the same plane; and vector C is given be to in another plane so their resultant never be zero except they are not perpendicular planes.
ankit singh
3 years ago
A + B + C = (a + c + g)i + (b + d + h)j + mk
Explanation:
Let the vectors A and B lie in the plane XY (arbitrary XY plane) and the vector is present outside the plane or any where in the space then
vectors A and B have two co ordinates that is
A = ai + bj
B = ci + dj
But vector C has three co ordinates that is
C = gi + hj + mk
where a , b , c, d ,g, ,h ,m are real numbers
and i , j , k are unit vectors along X ,Y and Z directions respectively.
NOW
A + B + C =  ai + bj + ci + dj + gi + hj + mk
A + B + C = (a + c + g)i + (b + d + h)j + mk
Hence
A + B + C = (a + c + g)i + (b + d + h)j + mk